OpenStudy (anonymous):

What is the coefficient in fron of the"x"^2 when (2x+3)^4 is expanded?

OpenStudy (anonymous):

How did you get this?

OpenStudy (anonymous):

do you just leave the 3 out of it?

OpenStudy (anonymous):

i used Pascal's triangle: (a + b)^2 = 1 2 1 (a + b)^3 = 1 3 3 1 (a + b)^4 = 1 4 6 4 1 the right side of the equal signs are just the coefficients of the binomials expanded.

OpenStudy (anonymous):

sorry, I misread your question. I thought your were looking for the coefficient of the leading term. you want the coefficient of the x^2 term right?

OpenStudy (anonymous):

Yes, the coefficient if front of x^2.

OpenStudy (anonymous):

ok, so... you see the 6 in the (a + b)^4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that's the term 6a^2b^2. since your a=2x, b=3 6*(2x)^2*(3)^2 = 6*4x^2*9 = 6*4*9*x^2 = 216x^2 your coefficient for the x^2 term is 216

OpenStudy (anonymous):

still faster than if you were to multiply the whole thing out...

OpenStudy (anonymous):

SO is that the answer because it says the coeficcient in front of x^2

OpenStudy (anonymous):

SO is that the answer because it says the coeficcient in front of x^2

OpenStudy (anonymous):

yes... http://www.wolframalpha.com/input/?i=expand+%282*x%2B3%29%5E4

OpenStudy (anonymous):

Okay thank you so much for your help!

OpenStudy (anonymous):

np